Consider three-points Gaussian quadrature. The three points on the x-axis and the associated weights are:

x1 =0, x2 = -"\sqrt{3/5}" , x3 = "\sqrt{3/5}" ; w1 = 8/9, w2 = w3 = 5/9

Apply this 3-points Gaussian quadrature to the polynomial

P(x) = 2x5-x4+x2-1



1
Expert's answer
2021-11-15T16:45:35-0500

"P(x)=2x^5-x^4+x^2-1"

"P(x_1)=P(0)=-1"

"P(x_2)=P\big(-\sqrt{\tfrac{3}{5}}\big)=-2\cdot \big(\tfrac{3}{5}\big)^2\cdot \sqrt{\tfrac{3}{5}}—\big(\tfrac{3}{5}\big)^2+\tfrac{3}{5}-1=-\tfrac{18}{25}\cdot \sqrt{\tfrac{3}{5}}-\tfrac{19}{25}"

"P(x_3)=P\big(\sqrt{\tfrac{3}{5}}\big)=2\cdot \big(\tfrac{3}{5}\big)^2\cdot \sqrt{\tfrac{3}{5}}—\big(\tfrac{3}{5}\big)^2+\tfrac{3}{5}-1=\tfrac{18}{25}\cdot \sqrt{\tfrac{3}{5}}-\tfrac{19}{25}"


"\int\limits_{-1}^1P(x)dx\approx\sum_{i=1}^3w_iP(x_i)=w_1P(x_1)+w_2P(x_2)+w_3P(x_3)=\tfrac{8}{9}\cdot (-1)+\tfrac{5}{9}(P(x_2)+P(x_3))=-\tfrac{8}{9}+\tfrac{5}{9}\cdot \big(-\tfrac{38}{25}\big)=-\tfrac{8}{9}-\tfrac{38}{45}=-\tfrac{40+38}{45}=-\tfrac{78}{45}=-\tfrac{26}{15}=-1\tfrac{11}{15}"


Answer: "\int\limits_{-1}^1P(x)dx\approx-1\tfrac{11}{15}" .